Circular No. 5841

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERSEID METEORS 1993 Reports from numerous observers in Europe and North America indi-

cate that, although there was no meteor storm, impressive meteor activity was observed from both continents, with a high percentage of bright fireballs and enduring trains; J. Rao, observing from the Mediterranean Sea, noted one fireball as bright as mag –10. S. Nakano indicated unusually weak activity in Japan up to Aug. 11.8 UT. Preliminary information compiled by M. Gyssens, International Meteor Organization, from observations made in Croatia, Czech Republic, Germany, Belgium, France and England, tentatively suggests a zenithal hourly rate ZHR ∼ 40 on Aug. 11.85, rising to perhaps 100 by Aug. 12.04, after which there was a more significant increase until Aug. 12.13, when the ZHR may well have surpassed 300. M. Kidger and others, observing from the Canary Islands, were indicating ZHR ∼ 350-400 during a 10-min interval around Aug. 12.13, dropping closer to 100 during Aug. 12.14-12.18. D. W. E. Green and others, observing under less satisfactory conditions in Massachusetts, independently reported a peak around Aug. 12.13, with significantly less activity during Aug. 12.17-12.21. In New Mexico, A. Hale remarked that he seemed to be seeing the end of a period of activity around Aug. 12.14, with a prevalence of bright, long-pathed meteors; as the altitude of the radiant increased and conditions improved he noted more normal activity until Aug. 12.37.

SUPERNOVA 1992bt IN NGC 3780 R. R. Treffers, B. Leibundgut and A. V. Filippenko, University of

California at Berkeley; and M. W. Richmond, Princeton University, report the discovery of a probable supernova in NGC 3780 (α = 11^{h}36^{m}38^{s}, δ = +56^{o}33′.0, equinox 1950.0), ∼ 22′′ west and 0′′.5 north of the galaxy’s nucleus. The object was found during the Leuschner Observatory Supernova Search, which uses an automated 0.76-m telescope equipped with the Lawrence Berkeley Laboratory CCD camera. It is visible at magnitude R ∼ 16 on 1992 Dec. 19 but was not automatically detected by the software because this was the first image of NGC 3780 obtained for the search. It is also visible at roughly the same brightness in images taken on Dec. 22 and 25, 1993 Jan. 4, 11, 25, and Mar. 5. By Mar. 31 the object seems to have faded ∼ 0.8 mag. It could not be detected to limiting mag 17.5 ± 0.3 in images obtained on Apr. 20 and later. The lightcurve suggests that the supernova was of type II (plateau), but confirmation with a late-time spectrum is desirable.

1993 August 12 Brian G. Marsden